# Harnessing quantum back-action for time-series processing

### v2

@@ IV.2 Considering a finite amount of resources
In the previous section, we analyzed the effect of making use of an indirect measurement-based approach to QRC in the ideal scenario of infinite resources. Here we consider the case in which each expectation value extracted from the reservoir is computed with a finite number of shots $N_{s}$. To take this into account, it has been shown that it is enough to add a Gaussian noise dependent on $N_{s}$ and $g$ to the previously computed expectation values [mujal2023time]. Namely, for one and two spins observables the following statistical uncertainties are considered:
@@
For the RSP, one simply considers the asymptotic regime of very large $g$. In this context, it is important to consider the inherent inefficiency of the RSP compared to the OLP, as discussed in [mujal2023time]. This analysis takes into account the time resources required for executing each protocol. With the RSP, the time required to run the QRC scales quadratically with respect to the dimension of the time series, whereas for the OLP, this dependence is linear. Therefore, by fixing the number of shots available for the OLP, we can compute the required time and subsequently determine the number of shots available for the RSP within that same time frame. Approximately, we find that the number of shots available for the RSP is proportional to the number available for the OLP divided by $K$, the size of the dataset (see Appendix A).
At this point, we can proceed to look at $P_{R}$, with the same logic as in Figure 2, but in a finite resources scenario. We now expect to see the combined effect of the back-action introduced by the OLP and of the higher efficiency of the OLP with respect to the RSP (Fig. 5). In this analysis, we fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$. Again the $P_{R}$ is probed in $33\times 40$ points in the ghg-h parameter space, and each point is the average over 50 different reservoir realizations.
@@
Figure 5: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom) in the scenario of finite resources. We fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$ and computed $N_{s}^{\text{RSP}}$ consequently for both tasks ($\approx 1.5\cdot 10^{3}$ for the forward prediction task and $\approx 3\cdot 10^{3}$ for the short-term memory task). The optimal parameters and corresponding $P_{R}$ values for each task are: forward prediction — RSP: $h^{\prime*}=15.118$; OLP: $g^{*}=0.486$, $h^{*}=30.703$, $P_{R}=1.158$; short-term memory — RSP: $h^{\prime*}=1.245$; OLP $g^{*}=0.355$ , $h^{*}=0.170$ , $P_{R}=1.416$.
@@
In this case, the same clear message emerges from the analysis of the $P_{R}$ in both tasks: in the finite resources scenario defined above, using the OLP with properly tuned $g$ and $h$ leads to higher performance with respect to the RSP. This conclusion is conveyed by the red regions that appear in the plots for $h\geq 10^{-1}$ and $g\leq 1$. In these regions, the $P_{R}$ is substantially greater than $1$, meaning that the OLP is performing better than the best-performing RSP. This, as pointed out in the previous sections, is a combined effect of both the OLP being more efficient in terms of resources and of the positive effect that the back-action brings to the performance of the reservoir in the optimal regime for the measurement strength.
@@ IV.3 Comparison with a classical feedback-based protocol
So far we have compared the OLP driven by an optimal back-action with the restarting protocol. This allowed us to highlight the advantage of exploiting the information disturbance tradeoff with respect to an ideal but inefficient scenario were measurements do not affect the system. More recently, alternative online schemes have been proposed where the reservoir is restarted at every step but complemented with a classical feedback layer that reconstructs the reservoir state informed by the previous expectation values [Kobayashi2024]. This approach is undoubtedly practical and efficient, yet it relies entirely on a classical memory, which limits the exploitation of genuinely quantum resources. By contrast, in our indirect-measurement OLP protocol, although less information is extracted per measurement, the reservoir retains quantum coherences that constitute a form of quantum memory. We therefore expect that, at comparable resource scaling, the optimized OLP can outperform protocols that rely only on classical memory like the feedback-driven protocol proposed in [Kobayashi2024].
@@ V Conclusions and Discussion
In this work, we analyzed the effects of indirect measurements on the QRC algorithm. Our findings reveal that, contrary to previous assumptions, the unwanted back-action from indirect measurements can positively influence the performance of QRC algorithms. By optimizing the performance ratio $P_{R}$, we highlighted that an informed choice of measurement strength and external fields in the OLP allows overperforming the standard RSP. The study is performed on several sub-tasks derived from two main benchmark tasks, widely used in the literature. We also show that an optimal parameter choice for the OLP can be crucial to obtaining good performances depending on specific sub-tasks. Moreover, we benchmark the optimal OLP against the feedback-driven protocol, a linear-scaling alternative that relies on classical memory and a unitary feedback layer. Our results demonstrate that the ability of the OLP to partially preserve coherences in the reservoir state provides an advantage that incoherent feedback alone cannot match.
As the field transitions to experimental implementations, our results highlight the necessity of identifying optimal protocols, especially in the finite resources scenario. In this context, we combine the previous knowledge on the better scaling of the OLP compared to the RSP with the ability to positively affect the reservoir dynamics by tuning the measurement strength. Together, these findings emphasize the OLP as a compelling candidate for near-term experimental demonstrations of QRC.This practical orientation is essential as experimental interest in QRC continues to grow [Yelin2024, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, PhysRevApplied.20.014051], providing a clear roadmap for researchers aiming to harness the capabilities of quantum systems for data processing tasks.
@@
A tutorial version of our implementation of the QRC algorithm utilized in this work, along with the necessary data to partially reproduce some of the figures, is available at [git_tuto]. The complete dataset can be provided upon request.
@@ Acknowledgements
We would like to express our gratitude to Rodrigo Martínez-Peña for their attentive reading and insightful comments on this manuscript. P.M. also thanks Rodrigo Martínez-Peña, Gian Luca Giorgi, Miguel C. Soriano, and Roberta Zambrini for their stimulating discussions at IFISC in Mallorca, which contributed significantly to the success of this project. G.F. would also like to thank Roberto Tricarico and Lisa Bombieri for the careful reading of the manuscript and the fruitful discussions. G.F. acknowledges support from a ”la Caixa” Foundation (ID 100010434) fellowship. The fellowship code is LCF/BQ/DI23/11990070. This project has received funding from MICIN, the European Union, NextGenerationEU (PRTR-C17.I1), the Government of Spain (Severo Ochoa CEX2019-000910-S and FUNQIP), the ERC AdG CERQUTE and the AXA Chair in Quantum Information Science, the ERC AdG NOQIA, MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI, QUANTERA MAQS PCI2019-111828-2, QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union.
Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project; European Union’s Horizon 2020 program under the Marie Sklodowska-Curie grant agreement No 847648; “La Caixa” Junior Leaders fellowships, La Caixa” Foundation (ID 100010434): CF/BQ/PR23/11980043.
@@ V.2 Independent parameter optimization per direction
In this appendix we explore the effect of maximizing the performance ratio $P_{R}$ independently for each observable direction. We do this analysis for the short-memory task. Concretely, we consider three separate scenarios in which only the observables along a single axis (either $x$, $y$, or $z$) are used to compute $P_{R}$, and we determine the optimal parameters $(g,h)$ that maximize $P_{R}$ for that direction following the same procedure described in the main text. This per-direction analysis give intuition about the contribution of each direction to the total performance.
As shown in Figure 8, the $x$ direction contributes less to $P_{R}$ than the other two directions, which are very similar both in behaviour and in optimal parameters.
Figure 8: Performance ratio analysis for the short-term memory task, computed independently for each measurement direction. Each panel shows the ghg-h map when only observables in a single direction are used to evaluate $P_{R}$. The optimal parameters and corresponding $P_{R}$ values for each direction are: ($x$). RSP: $h^{\prime*}=0.170$; OLP : $g^{*}=0.150$, $h^{*}=0.170$, $P_{R}=0.981$ ($y$). RSP: $h^{\prime*}=0.084$; OLP: $g^{*}=0.355$, $h^{*}=0.041$, $P_{R}=1.145$ ($z$). RSP: $h^{\prime*}=0.084$; OLP: $g^{*}=0.355$, $h^{*}=0.041$, $P_{R}=1.159$.
Next, we take the observables corresponding to each direction’s independently optimized parameters and build a combined dataset used to solve the task. In Figure 9 we compare the capacities of this combined dataset with those obtained from the joint (global) optimization reported in the main text. The combined, per-direction strategy yields a small increase in the performance ratio, $P_{R}=1.040$ (vs. $P_{R}=1.029$ for the global optimisation). Although this increase has little impact on the summed capacity reported in the main text, the independent optimisation redistributes capacity from larger delays (the "tail") to shorter delays (the "head"): capacities at small delay $\eta$ are improved while long-$\eta$ performance is slightly reduced. Thus, for applications that prioritise short-delay memory, the per-direction strategy can be advantageous. Finally, since the $x$ direction contributes relatively little, one may reasonably omit $x$ observables when resources are constrained and focus on the more informative $y$ and $z$ directions.

### v3

@@ IV.2 Considering a finite amount of resources
In the previous section, we analyzed the effect of making use of an indirect measurement-based approach to QRC in the ideal scenario of infinite resources. Here, we consider the case in which each expectation value extracted from the reservoir is computed with a finite number of shots $N_{s}$. To take this into account, it has been shown that it is enough to add a Gaussian noise dependent on $N_{s}$ and $g$ to the previously computed expectation values [mujal2023time]. Namely, for one and two spins observables the following statistical uncertainties are considered:
@@
For the RSP, one simply considers the asymptotic regime of very large $g$. In this context, it is important to consider the inherent inefficiency of the RSP compared to the OLP, as discussed in [mujal2023time]. This analysis takes into account the time resources required for executing each protocol. With the RSP, the time required to run the QRC scales quadratically with respect to the dimension of the time series, whereas for the OLP, this dependence is linear. Therefore, by fixing the number of shots available for the OLP, we can compute the required time and subsequently determine the number of shots available for the RSP within that same time frame. Approximately, we find that the number of shots available for the RSP is proportional to the number available for the OLP divided by $K$, the size of the dataset (see Appendix A). Finally, for completeness, we also consider the rewinding protocol (RWP). Since the RWP shares the same linear scaling in time as the OLP, both protocols can be allocated the same number of shots for a fair comparison at finite resources. This makes the RWP a natural unperturbed-dynamics benchmark in the finite-resource setting, allowing us to isolate the contribution of measurement back-action without the additional overhead present in the RSP.
As a first step, we focus on the practical and strict performance indicator $P_{R}^{\min}$ applied to the RSP. The goal is to quantify, in a realistic setting with finite resources, the combined effect of measurement back-action and the intrinsic scaling limitations of the RSP. Since the RSP becomes increasingly inefficient as the length of the time series grows, it provides a relevant baseline for evaluating how back-action interacts with sampling noise and resource constraints.
Using $P_{R}^{\min}$ is particularly appropriate in this context: by comparing the OLP with the best-performing unperturbed reservoir across all scanned parameters, this measure allows us to visualize the behaviour precisely in the regime where the reservoir performs well and where meaningful improvements are most significant. In other words, $P_{R}^{\min}$ highlights whether back-action can still provide an advantage even when the unperturbed dynamics is operating close to its optimal configuration, making it a stringent and practically motivated first benchmark before moving to more efficient protocols such as the RWP. We therefore compute $P_{R}^{\min}$ for the RSP (Fig. 6). In this analysis, we fix $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$. As before, $P_{R}^{\min}$ is evaluated on a $33\times 40$ grid in the gghh parameter space, and each point corresponds to an average over 50 different reservoir realizations.
@@
Figure 6: Minimum performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom) in the scenario of finite resources. We fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$ and computed $N_{s}^{\text{RSP}}$ consequently for both tasks ($\approx 1.5\cdot 10^{3}$ for the forward prediction task and $\approx 3\cdot 10^{3}$ for the short-term memory task). The optimal parameters and corresponding $P_{R}^{\min}$ values for each task are: forward prediction — RSP: $h^{\prime*}=15.118$; OLP: $g^{*}=0.486$, $h^{*}=30.703$, $P_{R}^{\min}=1.158$; short-term memory — RSP: $h^{\prime*}=1.245$; OLP $g^{*}=0.355$ , $h^{*}=0.170$ , $P_{R}^{\min}=1.416$.
@@
In this case, the same clear message emerges from the analysis of $P_{R}^{\min}$ in both tasks: in the finite-resources scenario defined above, using the OLP with properly tuned $g$ and $h$ leads to higher performance with respect to the unperturbed dynamics of the RSP, which is intrinsically inefficient at finite resources. This conclusion is conveyed by the red regions that appear in the plots for $h\geq 10^{-1}$ and $g\leq 1$. In these regions, $P_{R}^{\min}$ is substantially greater than $1$, meaning that the OLP is performing better than the best-performing RSP under the same resource budget. As discussed earlier, this behaviour reflects the combined effect of the higher efficiency of the OLP at finite resources and the positive contribution of back-action in the regime where the measurement strength is optimally tuned.
While this analysis is very useful in practice, it is difficult to isolate the true contribution of back-action when comparing the OLP with the RSP, since the finite-resource inefficiency of the RSP also plays an important role. For this reason, we complement the previous study with a comparison against the RWP. The RWP provides an efficient unperturbed-dynamics benchmark, as it shares the same linear scaling as the OLP; both protocols can therefore be allocated the same number of resources. This allows us to disentangle the effects of sampling noise from those of measurement back-action.
In this setting, we consider the maximal value of $P_{R}(g,h)$, now defined with respect to the RWP, ensuring a fair comparison between protocols with identical resource budgets. As shown in Fig. 7, we compute this quantity for different total numbers of resources and observe that the contribution of back-action remains significant even at finite sampling. Increasing the number of resources improves the performance, as the OLP requires higher sampling resolution to fully exploit the expectation values of its observables. This behaviour shows that the advantage observed in the ideal scenario is not an artifact of assuming infinite resources, but persists in practical conditions and progressively approaches the ideal limit as more measurements are available.
Figure 7: Maximum value of $P_{R}$ when comparing the OLP with the RWP, as a function of the total number of measurements $N_{s}$, for the forward prediction task (top) and for the short-term memory task (bottom). The red dashed line marks the threshold $P_{R}=1$. For both tasks, the maximal $P_{R}$ remains above one even at modest sampling, indicating that back-action provides a positive contribution to the performance of the OLP. As $N_{s}$ increases, the maximal $P_{R}$ grows and approaches the ideal values obtained in the infinite-resources scenario.
@@ IV.3 Comparison with a classical feedback-based protocol
So far we have compared the OLP driven by an optimal back-action with the restarting protocol. This allowed us to highlight the advantage of exploiting the information disturbance tradeoff with respect to an ideal but inefficient scenario were measurements do not affect the system. More recently, alternative online schemes have been proposed where the reservoir is restarted at every step but complemented with a classical feedback layer that reconstructs the reservoir state informed by the previous expectation values [Kobayashi2024, monomi2025feedbackenhancedquantumreservoircomputing]. This approach is undoubtedly practical and efficient, yet it relies entirely on a classical memory, which limits the exploitation of genuinely quantum resources. By contrast, in our indirect-measurement OLP protocol, although less information is extracted per measurement, the reservoir retains quantum coherences that constitute a form of quantum memory. We therefore expect that, at comparable resource scaling, the optimized OLP can outperform protocols that rely only on classical memory like the feedback-driven protocol proposed in [Kobayashi2024].
@@ V Conclusions and Discussion
In this work, we demonstrate that back-action, a purely quantum effect absent in classical systems, can be exploited as a computational resource, leading to enhanced performance of quantum reservoir computing in temporal data processing tasks.
By studying the performance ratio $P_{R}$, we showed that a suitable choice of the measurement strength and external field enables the OLP to match or even surpass the performance of unperturbed-dynamics protocols. In particular, large regions of the $(g,h)$ parameter space display a clear advantage for the OLP, with the strongest enhancements appearing precisely where the unperturbed dynamics is most constrained. To further characterize this effect, we introduced the stricter measure $P_{R}^{\min}$, which compares the OLP with the best-performing unperturbed reservoir. This allowed us to isolate the contribution of back-action independently of parameter choice. Notably, even for OLP configurations that already achieve high capacities, we observe that back-action continues to provide a measurable improvement, demonstrating that its beneficial role extends across the relevant operating regimes.
We then extended the analysis to finite-resource scenarios, where both sampling noise and protocol efficiency become relevant. When comparing the OLP with the RSP, the observed enhancements combine the intrinsic advantage provided by back-action with the superior scaling properties of the OLP. To disentangle these effects, we next compared the OLP with the RWP, which shares the same linear scaling in resources. By examining the maximal value of $P_{R}$ with respect to the RWP, we find that the positive impact of back-action remains significant even at finite sampling and converges toward the ideal infinite-resource behaviour as the number of shots increases. This confirms that the advantage observed for the OLP is a genuine effect of back-action rather than a consequence of differing resource requirements.
Across several sub-tasks derived from two standard benchmark problems, we also highlighted the importance of tuning the OLP parameters. The optimal measurement strength and external fields depend on the specific computational goal, and choosing these parameters appropriately is crucial to unlocking the benefits of back-action.
Finally, by benchmarking the OLP against the feedback-driven protocol, a linear-scaling classical–quantum hybrid scheme, we demonstrated that the OLP’s ability to partially preserve coherences in the reservoir state provides an advantage inaccessible to purely classical feedback mechanisms. Taken together, these findings establish that measurement back-action itself serves as a genuine computational resource for quantum reservoir computing.
As the field transitions to experimental implementations, our results highlight the importance of identifying protocols that make the best use of limited resources while exploiting the beneficial role of measurement back-action. In this context, we combine the previous knowledge on the better scaling of the OLP compared to the RSP with the ability to positively affect the reservoir dynamics by tuning the measurement strength. Together, these findings emphasize the OLP as a compelling candidate for near-term experimental demonstrations of QRC. This practical orientation is essential as experimental interest in QRC continues to grow [Yelin2024, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, PhysRevApplied.20.014051]. From an experimental perspective, the OLP is feasible with currently available technology. Its main requirement is the ability to implement weak or partial measurements with a tunable strength, a capability that has already been demonstrated in several quantum platforms, including superconducting circuits, cold atoms, and photonic systems [pan2020weak, spagnolo2022experimental, murch2013observing, naghiloo2019introduction, Yasuda2023]. In integrated photonics for example, such measurements can be naturally realized by using parametrized beamsplitters or tunable directional couplers that route the signal mode into an ancillary path only with a controlled probability. By detecting the ancilla mode while leaving the main mode largely undisturbed, these devices effectively implement a weak measurement whose strength is set by the beamsplitter transmissivity. This approach, together with the ability to incorporate additional ancilla modes on-chip, provides a practical route for realizing indirect-measurement-based protocols such as the OLP on scalable photonic hardware.
@@
A tutorial version of our implementation of the QRC algorithm utilized in this work, along with the necessary data to partially reproduce some of the figures, is available at [git_tuto]. The complete dataset is available at [data3058_2026].
@@ Acknowledgements
We would like to express our gratitude to Rodrigo Martínez-Peña for their attentive reading and insightful comments on this manuscript. P.M. also thanks Rodrigo Martínez-Peña, Gian Luca Giorgi, Miguel C. Soriano, and Roberta Zambrini for their stimulating discussions at IFISC in Mallorca, which contributed significantly to the success of this project. G.F. would also like to thank Roberto Tricarico and Lisa Bombieri for the careful reading of the manuscript and the fruitful discussions. G.F. acknowledges support from a ”la Caixa” Foundation (ID 100010434) fellowship. The fellowship code is LCF/BQ/DI23/11990070. This project has received funding from MICIN, the European Union, NextGenerationEU (PRTR-C17.I1), the Government of Spain (Severo Ochoa CEX2019-000910-S and FUNQIP), the ERC AdG CERQUTE and the AXA Chair in Quantum Information Science, the ERC AdG NOQIA, MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National STAMEENA PID2022-139099NB, the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI); Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; CEX2024-001490-S [MICIU/AEI/10.13039/501100011033]; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program), Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168268), EU Horizon Europe Program (No 101080086 NeQSTGrant Agreement 101080086 — NeQST).
@@ V.2 Independent parameter optimization per direction
In this appendix we explore the effect of maximizing the minimum performance ratio $P_{R}^{\min}$ independently for each observable direction. We do this analysis for the short-memory task. Concretely, we consider three separate scenarios in which only the observables along a single axis (either $x$, $y$, or $z$) are used to compute $P_{R}^{\min}$, and we determine the optimal parameters $(g,h)$ that maximize $P_{R}^{\min}$ for that direction following the same procedure described in the main text. This per-direction analysis give intuition about the contribution of each direction to the total performance.
As shown in Figure 10, the $x$ direction contributes less to $P_{R}^{\min}$ than the other two directions, which are very similar both in behaviour and in optimal parameters.
Figure 10: Minimum performance ratio analysis for the short-term memory task, computed independently for each measurement direction. Each panel shows the ghg-h map when only observables in a single direction are used to evaluate $P_{R}^{\min}$. The optimal parameters and corresponding $P_{R}^{\min}$ values for each direction are: ($x$). RSP: $h^{\prime*}=0.170$; OLP : $g^{*}=0.150$, $h^{*}=0.170$, $P_{R}^{\min}=0.981$ ($y$). RSP: $h^{\prime*}=0.084$; OLP: $g^{*}=0.355$, $h^{*}=0.041$, $P_{R}^{\min}=1.145$ ($z$). RSP: $h^{\prime*}=0.084$; OLP: $g^{*}=0.355$, $h^{*}=0.041$, $P_{R}^{\min}=1.159$.
Next, we take the observables corresponding to each direction’s independently optimized parameters and build a combined dataset used to solve the task. In Figure 11 we compare the capacities of this combined dataset with those obtained from the joint (global) optimization reported in the main text. The combined, per-direction strategy yields a small increase in the performance ratio, $P_{R}^{\min}=1.040$ (vs. $P_{R}^{\min}=1.029$ for the global optimisation). Although this increase has little impact on the summed capacity reported in the main text, the independent optimisation redistributes capacity from larger delays (the "tail") to shorter delays (the "head"): capacities at small delay $\eta$ are improved while long-$\eta$ performance is slightly reduced. Thus, for applications that prioritise short-delay memory, the per-direction strategy can be advantageous. Finally, since the $x$ direction contributes relatively little, one may reasonably omit $x$ observables when resources are constrained and focus on the more informative $y$ and $z$ directions.